The entrance to a tent is the shape of an equilateral triangle. If the base of the tent is 18 feet wide, then how tall is the tent?
Answer by stanbon(75887)

The entrance to a tent is the shape of an equilateral triangle. If the base of the tent is 18 feet wide, then how tall is the tent?
-------------------------
Draw the equilateral triangle.
Draw a perpendicular bisector from the top vertex to the base.
You now have two right triangles.
Use Pythagoras to solve for the height.
EQUATION:
h^2 + 9^2 = 18^2
h^2 = 18^2 - 9^2
h^2 = 243
height = sqrt243 = 15.59 feet
==============================
Cheers,
Stan H.